PHOTOELECTRON AND PHOTOION SPECTROSCOPY OF SCANDIUM-METHYLBENZENE COM-PLEXES

\underline{JUNG SUP LEE}, YUXIU LEI, SUDESH KUMARI, and DONG-SHENG YANG, Department of Chemistry, University of Kentucky, Lexington, KY 40506-0055.

We report here single UV photon pulsed field ionization-zero electron kinetic energy (PFI-ZEKE) photoelectron spectra of jet-cooled Sc-[C_{6}H_{6−n}(CH_{3})_{n}] (n = 2, 3, and 6) complexes and two-photon IR-UV photoionization spectra of Sc-[C_{6}(CH_{3})_{6}]. Adiabatic ionization energies and low-frequency metal-ligand and ligand-based vibrational modes are measured from the single-photon PFI-ZEKE spectra, and C-H stretching frequencies are determined from the two-photon IR-UV spectra. The ionization energies of these Sc complexes decrease with increasing methyl substations of hydrogen in the benzene ring. The C-H frequencies of hexamethylbenzene are significantly shifted due to Sc coordination. The ZEKE transitions of the di- and tri-methyl substituted complexes are sharp, while those of the hexa-methyl species are rather broad. The observed electronic transitions are ^{1}A ← ^{2}A for Sc-[1,4-C_{6}H_{4}(CH_{3})_{2}] (C_{2}), ^{1}A′ ← ^{2}A′ for Sc-[1,3,5-C_{6}H_{3}(CH_{3})_{3}] (C_{s}), and ^{1}A_{1} ←^{2} A_{1} for Sc-[C_{6}(CH_{3})_{6}] (C_{2v}).